For the best system, calculate the ratio of the masses of the buffer components required to make the buffer.

System is m(NH3)/m(NH4Cl)

Buffer has a pH=9.19.
Answer #1

Henderson's equation for Basic Buffer is

pH = pKa + log [ B ] / [ BH+]

In our case, pH = pKa + log [ NH3] / [ NH4Cl]

9.19 = 9.245 + log [ NH3] / [ NH4Cl]

log [ NH3] / [ NH4Cl] = 9.19-9.245 = - 0.055

[ NH3] / [ NH4Cl] = 10 - 0.055 = 0.881

i e   [ NH3] = 0.881 x [ NH4Cl]

We can write , No of moles of NH3 = 0.881 x No of moles of NH4Cl

Molar mass of NH3 = 14.01 + ( 3 x 1.0079) = 17.03 g / mol

Molar mass of NH4Cl = 14.01+ ( 4 x 1.0079) + 35.45 = 53.49 g / mol

Assume No of moles of NH3 = 1 mol then No of moles of NH4Cl = 1.14 mol.

Mass of 1 mol NH3 = 1 mol x ( 17.03 g / 1 mol ) = 17.03 g

Mass of 1.14 mol NH4Cl = 1.14 mol x ( 53.49 g / 1 mol) = 60.98 g

Therefore, Ratio of ( mass of  NH3 / mass of NH4Cl ) = 17.03 g / 60.98 g =0.2793